Adaptive Optics NIR Imaging of R136 in 30Doradus: The Stellar Population of a Nearby Starburst”

Bernhard Brandl

Max Planck Institut fuer extraterrestrische Physik Giessenbachstrasse

We report 0.′′15 resolution near infrared imaging of R136, the central region of 30 Doradus in the LMC. Our 12.′′8 × 12.′′8 images were recorded with the MPE camera SharpII at the 3.6m ESO telescope, using the adaptive optics system ComeOn+. The diffraction-limited spatial resolution and high sensitivity (20th magnitude in K) of our observations allowed our H and K band images to be compared and combined with recent HST WFPC2 data of R136. We were able to fit theoretical models with variable foreground extinction to the observed UVIHK magnitudes and thereby derive the stellar population in this starburst region. We find no red giants or supergiants, but ca. 100 extremely red sources which are probably young, pre-main-sequence low or intermediate mass stars. We obtained narrow band images to identify previously known and several new Wolf-Rayet stars by their HeII (2.189μm) and Brγ (2.166μm) emission lines. The presence of WR stars and absence of red supergiants narrow the cluster age to 3 − 5 Myrs, although the low ratio of WR to O stars in the central region would favour an age of ∼ 3.5 Myr. For the O–stars the core radius is found to be 0.1pc, but appears to decrease for larger stellar masses. The slope of the mass function is Γ = −1.6 on average, but steepens with increasing distance from the cluster center from Γ = −1.3 in the inner 0.4pc to Γ = −2.2 for stars more massive than m > 12M_{⊙} outside 0.8pc. Both, core radius and mass function variations, indicate mass segregation, possibly due to the cluster’s dynamical evolution.